OpenStudy (anonymous):

How would I graph an expression like |x^(2)-1| I know I would first convert it to a piece wise function where:

OpenStudy (anonymous):

WHOA 2 LAZY 2 ANSWER BOO HOO

OpenStudy (anonymous):

x^(2)-1 if x^(2) -1 >=0 and -x^(2) + 1 x^(2) -1 <0

OpenStudy (anonymous):

FIGURE IT OUTIT IS YOUR HW

OpenStudy (anonymous):

Isn't it just the graph of x^2 - 1 but with the part which is below the x-axis inverted so that it is above the x-axis? You will lose differentiability at x=-1 and x=1 but continuity is preserved so there is no need to define it piece-wise?

OpenStudy (anonymous):

MAKE ME

OpenStudy (anonymous):

So x^(2)-1 if x^(2) -1 >=0 and -x^(2) + 1 x^(2) -1 <0 x^(2) - 1 if x >= 1 and -x^(2) + 1 if x < 1 Thus I can just input values into both sides, anyone able to provide me with a more difficult piece wise function. I know it is elementary I just want to refresh my memory.

myininaya (myininaya):

@Hermella please don't yell at people to just figure it out on their own. They come on this site to get help. If you don't want to help, then just ignore their question.

myininaya (myininaya):

So how would you just graph x^2-1? this is a parabola right with y-intercept -1 and x-intercept -1 and 1

myininaya (myininaya):